OpenStudy (anonymous):

What fraction is between 2 and 3

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

thanks all

OpenStudy (anonymous):

yep. :)

OpenStudy (anonymous):

No problemo. :D

OpenStudy (anonymous):

Any fraction with a value between 2 and 3 2 and (1/2) = (5/2) 2 and 1 tenth = (21/10)

OpenStudy (mathstudent55):

There is an infinite number of fractions whose value is between 2 and 3. Anyone will do? For example, 21/10, which is the same as 2.1 is between 2 and 3.

OpenStudy (anonymous):

33/100 is also an option

OpenStudy (anonymous):

(1/2) is not between 2 and 3. (1/2) is between 0 and 1

OpenStudy (anonymous):

Several integers are between including decimal points

OpenStudy (mathstudent55):

33/100 = 0.33, which is less than 2, so it's not between 2 and 3.

OpenStudy (anonymous):

Potatoes

OpenStudy (anonymous):

No I did not say 2

OpenStudy (anonymous):

@mattdogg What? how is that relevant

OpenStudy (mathstudent55):

There is no integer between 2 and 3 since 2 and 3 are consecutive integers.

OpenStudy (anonymous):

I calculated by

OpenStudy (anonymous):

I am done.

OpenStudy (anonymous):

integers are whole numbers

OpenStudy (anonymous):

Not all

OpenStudy (anonymous):

Sorry auto correct

OpenStudy (mathstudent55):

Natural numbers: 1, 2, 3, 4, 5, 6, ... Whole numbers: 0, 1, 2, 3, 4, 5, 6, ... Integers: ..., -3, -2, -1, 0, 1, 2, 3, ...